OpenStudy (anonymous):

ANYONE!! ??Prehistoric cave paintings were discovered in a cave in France. The paint contained 14% of the original Carbon-14. Estimate the age of the paintings.

OpenStudy (unklerhaukus):

So you know the half life \(t_{1/2}\) of \(^{14}C\) is \(5730 \text {yr}\)

OpenStudy (unklerhaukus):

\[N(t)=N_02^{-t/t_{1/2}}\] \[\frac{N(t)}{N_0}=2^{-t/t_{1/2}}\]

OpenStudy (unklerhaukus):

\[\frac{N(t)}{N_0}=\frac{14\%}{100\%}=14\%\]

OpenStudy (unklerhaukus):

take long to base two